# A Low-Overhead Incorporation-Extrapolation based Few-Shot CSI Feedback Framework for Massive MIMO Systems

### v1

@@ III-C Frequency Extrapolation Network
In contrast, to extrapolate $\hat{\overline{\mathbf{W}}}$ to $\hat{\mathbf{W}}$ with low computational complexity, we design a lightweight but effective frequency extrapolation network by exploiting multi-path extrapolation and residual learning based on the subcarrier-level and the group-level frequency domain correlations, as shown in Fig. 1. The frequency extrapolation network is composed of $N_{gr}$ extrapolation modules (EMs) and one concatenation module. Each EM further consists of two linear layers, one Gaussian Error Linear Unit (GELU) activation layer, one residual connection operation[35], and one layer normalization operation[36]. The residual connection is designed to allow the activation from one layer to be directly passed to a later layer and thus bypass the intermediate layers. This avoids the issue of vanishing gradients in EMs and enables designing deep neural network architectures with improved model capacity for frequency domain extrapolation. Specifically, denoting the input to the frequency extrapolation network as $\mathbf{X}^{de}\in\mathbb{R}^{(2\times N_{T})\times N_{grp}}$, $\mathbf{X}^{de}$ is first processed by $EM_{i}$, $i=1,2,\ldots,N_{gr}$, as
@@
where $\mathbf{X}^{EM_{i}}(1)$, $\mathbf{X}^{EM_{i}}(1)$, and $\mathbf{X}^{EM_{i}}$ are the outputs of the GELU activation layer, the residual connection operation, and the layer normalization operation of the $i$-th EM, respectively. $\operatorname{GELU}(\cdot)$ denotes the GELU activation function, $\mathbf{K}_{1}^{i}\in\mathbb{R}^{N_{grp}\times N_{grp}}$ and $\mathbf{K}_{2}^{i}\in\mathbb{R}^{N_{grp}\times N_{grp}}$ are learnable weight matrices, $\mathbf{b}_{1}^{i}\in\mathbb{R}^{1\times N_{grp}}$ and $\mathbf{b}_{2}^{i}\in\mathbb{R}^{1\times N_{grp}}$ are learnable bias vectors, $\mathbb{E}_{-1}\{\cdot\}$ and $\operatorname{Var}_{-1}\{\cdot\}$ are the expectation and variance of the input matrix along its last dimension, $\odot$ denotes the Hadamard product, $\mathbf{g}_{i}\in\mathbb{R}^{1\times N_{grp}}$ and $\mathbf{l}_{i}\in\mathbb{R}^{1\times N_{grp}}$ are learnable affine transformation parameters, and $\epsilon$ is a small real number (e.g. 1e-5) added to the denominator for numerical stability. Then, the outputs of these $N_{gr}$ EMs are concatenated to form the final output $\mathbf{X}^{ext}\in\mathbb{R}^{(2\times N_{T})\times N_{c}}$ as
@@
Hence, $\hat{\mathbf{W}}$ can be achieved from $\mathbf{X}^{ext}$ by first reshaping $\mathbf{X}^{ext}$ to $\dot{\mathbf{X}}^{ext}\in\mathbb{R}^{2\times N_{T}\times N_{c}}$ and then splitting $\dot{\mathbf{X}}^{ext}$ along its first dimension to obtain the real part and the imaginary part of $\hat{\mathbf{W}}$.
It is worth emphasizing that in the frequency extrapolation network, each EM mainly exploits the group-level frequency domain correlation while the multi-path learning architecture mainly exploits the subcarrier-level frequency domain correlation. And we also emphasize that the frequency extrapolation network is effective but lightweight for the BS side to handle the extrapolation process. This is because the proposed frequency extrapolation network performs frequency domain extrapolation at the group level owing to the multi-path architecture, and thus avoids direct processing at the subcarrier level and achieves significant computational efficiency since $N_{grp}\ll N_{c}$.
@@ III-D Model Training
In the model training stage, the proposed Transformer-based CSI feedback network and the proposed frequency extrapolation network are trained jointly in an end-to-end manner. The loss function used for model training is defined as
@@
$\rho_{i}=\frac{\|\mathbf{w}_{i}^{H}\hat{\mathbf{w}}_{i}\|_{2}}{\|\mathbf{w}_{i}\|\|_{2}\hat{\mathbf{w}}_{i}\|_{2}},$ (22)
@@
where $\mathbb{E}(\cdot)$ is the expectation over a batch of training samples, $F_{en}$ and $F_{de}$ represent the encoder and decoder of the Transformer-based CSI feedback network with learnable parameters $\mathbf{\Theta}_{en}$ and $\mathbf{\Theta}_{de}$, respectively, and $F_{ext}$ represents the frequency extrapolation network with learnable parameters $\mathbf{\Theta}_{ext}$.
@@ IV-A Knowledge-Driven Data Augmentation Method
According to the channel property, there are both subcarrier-level and group-level frequency domain correlations that can be leveraged to augment data samples. Based on this domain knowledge, we propose the knowledge-driven data augmentation method for CSI feedback. Recall that in the incorporation process, every $N_{gr}$ subcarriers in $\mathcal{H}_{subc}$ are grouped together and each group $\mathcal{H}_{grp}^{j}$ yields one dominant eigenvector $\overline{\mathbf{w}}_{j}$. Intuitively, $\overline{\mathbf{W}}=[\overline{\mathbf{w}}_{1},\overline{\mathbf{w}}_{2},\ldots,\overline{\mathbf{w}}_{N_{grp}}]$ corresponding to all groups $\{\mathcal{H}_{grp}^{j}|j=1,2,\ldots,N_{grp}\}$ from collected samples can be directly used for training the CSI feedback network. To augment the training samples by fully leveraging the frequency domain correlations, we also consider a small granularity $N_{sgr}<N_{gr}$. [^6] To keep high frequency domain correlations in the augmented samples, we only consider a small granularity, i.e., $N_{sgr}<N_{gr}$. As a result, the $j$-th group, i.e., $\mathcal{H}_{grp}^{j}$, can be further divided into $N_{sgrp}=N_{gr}/N_{sgr}$ subgroups, i.e., $\mathcal{H}_{sgrp}^{j,k}$, $j=1,2,\ldots,N_{grp}$ and $k=1,2,\ldots,N_{sgrp}$. And we have
@@
where $\mathbf{m}=[m_{1},m_{2},\ldots,m_{N_{grp}}]$ and $m_{n}\in\{1,2,\ldots,N_{sgrp}\}$ for $n=1,2,\ldots,N_{grp}$. Therefore, theoretically, with different settings of $\mathbf{m}$, ${(C_{N_{sgrp}}^{1})}^{N_{grp}}$ (here $C_{a}^{b}$ denotes the combination operator indicating choosing $b$ elements from $a$ elements) augmented samples can be generated for $\overline{\mathbf{W}}$. For ease of understanding, Fig. 2 gives an illustration of the knowledge-driven data augmentation method with $N_{gr}=4$ and $N_{sgr}=2$, where $N_{grp}$ eigenvectors (marked with red ticks) from subgroups are selected to form an augmented eigenvector-based CSI matrix $\widetilde{\mathbf{W}}_{[1,2,\ldots,1]}$ with $\mathbf{m}=[1,2,\ldots,1]$.
Note that the proposed knowledge-driven data augmentation method can be easily implemented to yield extensive augmented samples, which is very efficient in improving the performance of CSI feedback, especially under extremely few shot CSI feedback scenarios. Nevertheless, such augmented samples are generated relying on frequency domain correlations from limited number of practically collected samples, which thus lack of sample diversity. To further improve the diversity of augmented samples, we resort to AIGC-based data augmentation methods and propose the EGAN to generate diverse samples by learning the distribution of data samples.
@@ IV-B AIGC-based Data Augmentation Method
Existing AIGC-based data augmentation methods for few-shot CSI feedback augment data samples by first learning the distribution of collected data samples and then generating diverse samples following the learned distribution via generative models (e.g., GANs). However, compared with discriminative models, generative models typically require large amounts of training data[31] to model the distribution well, which contradicts the intention of exploiting AIGC for data augmentation under few-shot scenarios. To solve this issue, we propose a novel AIGC-based data augmentation method, which is able to model the distribution of the CSI matrices well with a few collected samples. Specifically, a new GAN architecture, called EGAN, is proposed to learn the distributions of CSI matrices on two scales, i.e., the low-dimensional CSI matrices $\overline{\mathbf{W}}$ and the full-dimensional CSI matrices $\mathbf{W}$. Compared with vanilla GANs, the EGAN benefits from learning an additional distribution of low-dimensional CSI matrices, given that low-dimensional CSI matrices contain internal information of full-dimensional CSI matrices thanks to the incorporation process. This is beneficial for the generation of full-dimensional CSI matrices. Moreover, the knowledge-driven data augmentation method is also adopted to generate augmented data samples for training the EGAN with sufficient samples under few-shot scenarios.
@@ IV-B2 Training of EGAN
$+\lambda_{1}\mathbb{E}_{\tilde{\underline{\mathbf{X}}}^{GL}\sim p^{\tilde{GL}}}\{D_{1}(\underline{\mathbf{X}}^{GL})\}\left.+\lambda_{2}\mathbb{E}_{\tilde{\underline{\mathbf{X}}}^{GF}\sim p^{\tilde{GF}}}\{D_{2}(\underline{\mathbf{X}}^{GF})\}\right),$ (41)
where $\tilde{\underline{\mathbf{X}}}^{GL}=\alpha\underline{\mathbf{X}}^{GL}+(1-\alpha)\underline{\mathbf{X}}^{GL}$ and $\tilde{\underline{\mathbf{X}}}^{GF}=\alpha\underline{\mathbf{X}}^{GF}+(1-\alpha)\underline{\mathbf{X}}^{GF}$ are randomly interpolated samples with $\alpha\sim U[0,1)$, $U$ denotes the uniform distribution, and $\lambda_{1}$ and $\lambda_{2}$ are constant penalty coefficients.
Note that the training of EGAN is a min-max optimization problem and is solved by alternately optimizing the generator and two discriminators. Additionally, the knowledge-driven data augmentation method is adopted to generate augmented data samples as real low-dimensional samples to train the EGAN.
@@ V Simulation Results
In this section, we first introduce the experimental settings and the performance evaluation metrics. Then, we demonstrate the performance of the proposed two data augmentation methods in enabling CSI feedback under few-shot scenarios. Finally, we compare the proposed CSI feedback framework with existing schemes under different numbers of collected samples and feedback bits.
@@ V-A1 Experimental Datasets
In both scenarios, the BS is equipped with a uniform linear array (ULA) with $32$ antennas and the UE is equipped with a ULA with $4$ antennas. The systems both operate with OFDM modulation with $1024$ subcarriers. As a result, each channel sample is with the size of $4\times 32\times 1024$ and the corresponding full eigenvector-based CSI matrix is with the size of $32\times 1024$. The group granularity of the proposed framework is set as $N_{gr}=16$. Therefore, the size of the CSI matrix to be fed back is $32\times 64$.

### v2

@@ III-C Frequency Extrapolation Network
In contrast, to extrapolate $\hat{\overline{\mathbf{W}}}$ to $\hat{\mathbf{W}}$ with low computational complexity, we design a lightweight but effective frequency extrapolation network by exploiting multi-path extrapolation and residual learning based on the subcarrier-level and the group-level frequency domain correlations, as shown in Fig. 1. The frequency extrapolation network is composed of $N_{gr}$ extrapolation modules (EMs) and one concatenation module. Each EM further consists of two linear layers, one Gaussian Error Linear Unit (GELU) activation layer, one residual connection operation[38], and one layer normalization operation[39]. The residual connection is designed to allow the activation from one layer to be directly passed to a later layer and thus bypass the intermediate layers. This avoids the issue of vanishing gradients in EMs and enables the designing of deep neural network architectures with improved model capacity for frequency domain extrapolation. Specifically, denoting the input to the frequency extrapolation network as $\mathbf{X}^{de}\in\mathbb{R}^{(2\times N_{T})\times N_{grp}}$, $\mathbf{X}^{de}$ is first processed by $EM_{i}$, $i=1,2,\ldots,N_{gr}$, as
@@
where $\mathbf{X}^{EM_{i}}(1)$, $\mathbf{X}^{EM_{i}}(2)$, and $\mathbf{X}^{EM_{i}}$ are the outputs of the GELU activation layer, the residual connection operation, and the layer normalization operation of the $i$-th EM, respectively. $\operatorname{GELU}(\cdot)$ denotes the GELU activation function, $\mathbf{K}_{1}^{i}\in\mathbb{R}^{N_{grp}\times N_{grp}}$ and $\mathbf{K}_{2}^{i}\in\mathbb{R}^{N_{grp}\times N_{grp}}$ are learnable weight matrices, $\mathbf{b}_{1}^{i}\in\mathbb{R}^{1\times N_{grp}}$ and $\mathbf{b}_{2}^{i}\in\mathbb{R}^{1\times N_{grp}}$ are learnable bias vectors, $\mathbb{E}_{-1}\{\cdot\}$ and $\operatorname{Var}_{-1}\{\cdot\}$ are the expectation and variance of the input matrix along its last dimension, $\odot$ denotes the Hadamard product, $\mathbf{g}_{i}\in\mathbb{R}^{1\times N_{grp}}$ and $\mathbf{l}_{i}\in\mathbb{R}^{1\times N_{grp}}$ are learnable affine transformation parameters, and $\epsilon$ is a small real number (e.g. 1e-5) added to the denominator for numerical stability. Then, the outputs of these $N_{gr}$ EMs are concatenated to form the final output $\mathbf{X}^{ext}\in\mathbb{R}^{(2\times N_{T})\times N_{c}}$:
@@
Hence, $\hat{\mathbf{W}}$ can be achieved from $\mathbf{X}^{ext}$ by first reshaping $\mathbf{X}^{ext}$ to $\underline{\mathbf{X}}^{ext}\in\mathbb{R}^{2\times N_{T}\times N_{c}}$ and then splitting $\underline{\mathbf{X}}^{ext}$ along its first dimension to obtain the real part and the imaginary part of $\hat{\mathbf{W}}$.
It is worth emphasizing that in the frequency extrapolation network, each EM mainly exploits the group-level frequency domain correlation resorting to residual learning while the multi-path learning architecture mainly exploits the subcarrier-level frequency domain correlation to refine the concatenated outputs of all EMs. And we also emphasize that the frequency extrapolation network is effective but lightweight for the BS side to handle the extrapolation process. This is because the proposed frequency extrapolation network performs frequency domain extrapolation via multi-path learning and thus avoids direct extrapolation at the subcarrier level and achieves significant computational efficiency compared with existing convolutional extrapolation networks[40, 41, 42]. The computational complexity of the proposed frequency extrapolation network (FEN) and the convolutional extrapolation network (CEN) in [40, 41, 42] can be expressed by [^6] Note that we also consider a two-layer architecture similar to our frequency extrapolation network for the convolutional extrapolation network for a fair comparison.
$\mathcal{O}(FEN)$ $=\mathcal{O}(4N_{gr}N_{T}N_{grp}^{2})=\mathcal{O}(4N_{T}N_{grp}N_{c}),$ (21)
$\mathcal{O}(CEN)$ $=\mathcal{O}(2KN_{T}N_{grp}N_{c}+2KN_{T}N_{c}^{2}),$ (22)
where $K>=1$ is the kernel size of the convolutional extrapolation network. Referring to (21) and (22), the proposed extrapolation network is much more computationally efficient than the convolutional extrapolation network since $N_{grp}\ll N_{c}$.
@@ III-D Model Training
In the model training stage, the proposed Transformer-based CSI compression and reconstruction network and the proposed frequency extrapolation network are trained jointly in an end-to-end manner. The loss function used for model training is defined as
@@
$\rho_{i}=\frac{\|\mathbf{w}_{i}^{H}\hat{\mathbf{w}}_{i}\|_{2}}{\|\mathbf{w}_{i}\|_{2}\|\hat{\mathbf{w}}_{i}\|_{2}},$ (24)
@@
where $\mathbb{E}(\cdot)$ is the expectation over a batch of training samples, $F_{en}$ and $F_{de}$ represent the encoder and decoder of the Transformer-based CSI compression and reconstruction network with learnable parameters $\mathbf{\Theta}_{en}$ and $\mathbf{\Theta}_{de}$, respectively, and $F_{ext}$ represents the frequency extrapolation network with learnable parameters $\mathbf{\Theta}_{ext}$.
@@ IV-A Knowledge-Driven Data Augmentation Method
According to the channel property, there are both subcarrier-level and group-level frequency domain correlations that can be leveraged to augment data samples. Based on this domain knowledge, we propose the knowledge-driven data augmentation method for CSI feedback. Recall that in the incorporation process, every $N_{gr}$ subcarriers in $\mathcal{H}_{subc}$ are grouped together and each group $\mathcal{H}_{grp}^{j}$ yields one dominant eigenvector $\overline{\mathbf{w}}_{j}$. Intuitively, $\overline{\mathbf{W}}=[\overline{\mathbf{w}}_{1},\overline{\mathbf{w}}_{2},\ldots,\overline{\mathbf{w}}_{N_{grp}}]$ corresponding to all groups $\{\mathcal{H}_{grp}^{j}|j=1,2,\ldots,N_{grp}\}$ from collected samples can be directly used for training the CSI feedback network. To augment the training samples by fully leveraging the frequency domain correlations, we also consider a small granularity $N_{sgr}<N_{gr}$. [^7] To keep highly correlated frequency domain correlations in the augmented samples, we only consider a small granularity, i.e., $N_{sgr}<N_{gr}$. As a result, the $j$-th group, i.e., $\mathcal{H}_{grp}^{j}$, can be further divided into $N_{sgrp}=N_{gr}/N_{sgr}$ subgroups, i.e., $\mathcal{H}_{sgrp}^{j,k}$, $k=1,2,\ldots,N_{sgrp}$. And we have
@@
where $\mathbf{m}=[m_{1},m_{2},\ldots,m_{N_{grp}}]$ and $m_{n}\in\{1,2,\ldots,N_{sgrp}\}$ for $n=1,2,\ldots,N_{grp}$. Therefore, theoretically, with different settings of $\mathbf{m}$, ${(C_{N_{sgrp}}^{1})}^{N_{grp}}$ (here $C_{a}^{b}$ denotes the combination operator indicating choosing $b$ elements from $a$ elements) augmented samples can be generated for $\overline{\mathbf{W}}$. [^8] It is worth mentioning that we can also group non-continuous subcarriers together to further increase the number of augmented samples. However, grouping non-continuous subcarriers together would increase the computational complexity of the KDDA method but generate similar augmented samples, which is not very efficient. For ease of understanding, Fig. 2 gives an illustration of the knowledge-driven data augmentation method with $N_{gr}=4$ and $N_{sgr}=2$, where $N_{grp}$ eigenvectors (marked with red ticks) from subgroups are selected to form an augmented eigenvector-based CSI matrix $\widetilde{\mathbf{W}}_{[1,2,\ldots,1]}$ with $\mathbf{m}=[1,2,\ldots,1]$.
Note that the proposed knowledge-driven data augmentation method can be easily implemented to yield extensive augmented samples, which is very efficient in improving the performance of CSI feedback, especially under extremely few-shot CSI feedback scenarios. Nevertheless, such augmented samples are generated relying on frequency domain correlations from a limited number of practically collected samples, which thus lack of sample diversity. To further improve the diversity of augmented samples, we resort to AIGC-based data augmentation methods and propose the EGAN to generate diverse samples by learning the distribution of data samples.
@@ IV-B AIGC-based Data Augmentation Method
Existing AIGC-based data augmentation methods for few-shot CSI feedback augment data samples by first learning the distribution of collected data samples and then generating diverse samples following the learned distribution via generative models (e.g., GANs). However, compared with discriminative models, generative models typically require large amounts of training data[34] to model the distribution well, which contradicts the intention of exploiting AIGC for data augmentation under few-shot scenarios. To solve this issue, we propose a novel AIGC-based data augmentation method, which is able to model the distribution of the CSI matrices well with a few collected samples. Specifically, a new GAN architecture, called EGAN, is proposed to learn the distributions of CSI matrices on two scales, i.e., the low-dimensional CSI matrices $\overline{\mathbf{W}}$ and the full-dimensional CSI matrices $\mathbf{W}$. Compared with vanilla GANs, the EGAN benefits from learning an additional distribution of low-dimensional CSI matrices, given that low-dimensional CSI matrices contain internal information of full-dimensional CSI matrices thanks to the incorporation process. This is beneficial for the generation of full-dimensional CSI matrices. Moreover, as depicted in Fig. 3, the knowledge-driven data augmentation method is also adopted to generate augmented data samples for training the EGAN with sufficient samples under few-shot scenarios.
Figure 3: The diagram depicting the connection between the knowledge-driven data augmentation method and the AIGC-based data augmentation method.
@@ IV-B1 Network Architecture of the EGAN
Note that the outputs of the two discriminators, i.e., $\underline{\mathbf{X}}^{{D1,DS_{N_{D1}}}}$ and $\underline{\mathbf{X}}^{{D2,DS_{N_{D2}}}}$, are three-dimensional tensors as shown in Fig. 4, which can be regarded as distributed representations for the distributions of low-dimensional and full-dimensional samples, respectively. Compared with the vanilla GAN which outputs a probability indicating whether the input is a generated sample or a real sample, a distributed representation in a tensor form keeps more information about the sample distribution and is thus more effective[47].
@@ IV-B2 Training of EGAN
$+\lambda_{1}\mathbb{E}_{\tilde{\underline{\mathbf{X}}}^{L}\sim\tilde{p}^{L}}\{(\|\nabla_{\tilde{\underline{\mathbf{X}}}^{L}}D_{1}(\tilde{\underline{\mathbf{X}}}^{L})\|_{2}-1)^{2}\}$
$\left.+\lambda_{2}\mathbb{E}_{\tilde{\underline{\mathbf{X}}}^{F}\sim\tilde{p}^{F}}\{(\|\nabla_{\tilde{\underline{\mathbf{X}}}^{F}}D_{2}(\tilde{\underline{\mathbf{X}}}^{F})\|_{2}-1)^{2}\}\right),$ (43)
where $\tilde{\underline{\mathbf{X}}}^{L}\sim\tilde{p}^{L}=\alpha\underline{\mathbf{X}}^{L}+(1-\alpha)\underline{\mathbf{X}}^{GL}$ and $\tilde{\underline{\mathbf{X}}}^{F}\sim\tilde{p}^{F}=\alpha\underline{\mathbf{X}}^{F}+(1-\alpha)\underline{\mathbf{X}}^{GF}$ are randomly interpolated samples with $\alpha\sim U[0,1)$ for gradient penalty, $U$ denotes the uniform distribution, and $\lambda_{1}$ and $\lambda_{2}$ are constant penalty coefficients.
Note that the training of EGAN is a min-max optimization problem and is solved by alternately optimizing the generator and two discriminators. Specifically, the maximization problem (i.e., optimizing the two discriminators while the generator is fixed) and the minimization problem (i.e., optimizing the generator while the discriminators are fixed) are solved alternatively in an adversarial training manner. The discriminators are responsible for distinguishing real and generated samples and thus are optimized to maximize the Wasserstein distance, and the generator is optimized to minimize the Wasserstein distance for generating samples that are more real from the distribution perspective. Therefore, through the adversarial training, the Wasserstein distance will decrease gradually and finally, the generator will generate samples following a distribution similar to that of real samples.Additionally, the knowledge-driven data augmentation method is adopted to generate augmented data samples for training the EGAN as depicted in Fig. 3, such that the EGAN can be trained with few collected samples.
@@ V Simulation Results
In this section, we first introduce the experimental settings and the performance evaluation metrics. Then, we demonstrate the performance of the proposed two data augmentation methods in enabling CSI feedback under few-shot scenarios. After that, we compare the performance of the proposed CSI feedback framework with existing schemes under different numbers of collected samples, feedback bits, and signal-to-noise ratios (SNRs). Finally, we analyze the computational complexity of the proposed CSI feedback framework.
@@ V-A1 Experimental Datasets
In both scenarios, the BS is equipped with a uniform linear array (ULA) with $32$ antennas and the UE is equipped with a ULA with $4$ antennas. The systems both operate with OFDM modulation with $1024$ subcarriers. As a result, each channel sample is with the size of $4\times 32\times 1024$ and the corresponding full eigenvector-based CSI matrix is with the size of $32\times 1024$.
@@
The proposed incorporation-extrapolation CSI feedback scheme and the knowledge-driven data augmentation method essentially rely on frequency domain correlations. To quantify the frequency domain correlation among the the $i$-th subcarrier group and $(i+1)$-th subcarrier group of the $j$-th CSI sample given a group granularity $N_{gr}$, the adjacent GCS, denoted by $\tilde{\rho}_{j}(i,i+1)$, can be exploited, which is defined as
$\tilde{\rho}_{j}(i,i+1)=\frac{\|\mathbf{w}_{i}^{H}\mathbf{w}_{i+1}\|_{2}}{\|\mathbf{w}_{i}\|_{2}\|\mathbf{w}_{i+1}\|_{2}},$ (44)
where $\|\cdot\|_{2}$ denotes the L2 norm, and $\mathbf{w}_{i}$ and $\mathbf{w}_{i+1}$ are the dominant eigenvectors of the $i$-th and $i+1$-th subcarrier groups, respectively. Note that the closer the $\tilde{\rho}$ is to $1$, the more correlated the adjacent subcarrier groups are, while the closer the $\tilde{\rho}$ is to 00, the less correlated the adjacent subcarrier groups are. Then, the average frequency domain correlations over all adjacent subcarrier groups and collected samples, denoted by $\overline{\tilde{\rho}}$, can be defined as
$\overline{\tilde{\rho}}=\frac{1}{N_{col}(N_{grp}-1)}\sum_{j=1}^{N_{col}}\sum_{i=1}^{N_{grp}-1}\tilde{\rho}_{j}(i,i+1),$ (45)
where $N_{col}$ is the number of collected samples, $N_{grp}$ is the number of subcarrier groups.
Table III shows the average frequency domain correlations among adjacent subcarrier groups over all collected samples with different group granularities under the indoor and outdoor scenarios. According to Table III, to retain high frequency domain correlations (e.g., $\rho>0.95$) for CSI feedback and knowledge-driven data augmentation, the group granularity of the proposed framework is set as $N_{gr}=16$. Therefore, the size of the CSI matrix to be fed back is $32\times 64$.